In India during the 1980s the incremental capital-output ratio (ICOR) averaged 2.50 and



population growth rate and depreciation respectively, n=0.02, δ=0.01.



a. Using the Harrod growth equation, what saving rate would have been required for India to



achieve an aggregate growth rate of 8 percent per annum?



b. With the same ICOR, what growth target could be achieved with a saving rate of 27 percent?



c. If there is a large increase in the saving rate, and therefore a large increase in the amount of



new capital formation, is the ICOR likely to rise, fall, or remain the same? Explain.

1
Expert's answer
2022-01-25T08:45:39-0500

a) Growth rate

"g=\frac{s}{k}"

"s=g\times k"

=8"\times 2.5= 20%"%


b) "g=\frac{s}{k}"

Growth Target to be achieved by 27% savings rate with the same ICOR

"g=\frac{27}{2.5}= 10.8" %


c) It will rise because it is calculated as the growth in the capital stock divided by the growth in GDP. Since Investment is defined as the growth in the capital stock, the ICOR is equal to Investment divided by the growth of GDP. Since investment increases, then this translates to an increase in ICOR.